2x + 2y=8 and 4x-6y=36

solution (6,-2) how did my text book come up with thesesolutions

Answer by stanbon(75887)

2x + 2y=8
4x - 6y=36
-----------
Divide thru the 2nd equation by 2 to get:
2x + 2y = 8
2x - 3y = 18
-----------------
subtract the 2nd from the 1st equation to solve for "y":
5y = -10
y = -2
--------------
Substitute into 2x + 2y = 8 to solve for "x":
2x + 2*-2 = 8
2x - 4 = 8
2x = 12
x = 6
--------------------------------------------
solution (6,-2)
=============================================
Cheers,
Stan H.